# When Should We Introduce Safety Interventions During Pretraining?

### v1

@@ Abstract
Ensuring the safety of language models in high-stakes settings remains a pressing challenge, as aligned behaviors are often brittle and easily undone by adversarial pressure or downstream finetuning. Prior work has shown that interventions applied during pretraining, such as rephrasing harmful content, can substantially improve the safety of the resulting models. In this paper, we study the fundamental question: “When during pretraining should safety interventions be introduced?” We keep the underlying data fixed and vary only the choice of a safety curriculum: the timing of these interventions, i.e., after 0%, 20%, or 60% of the pretraining token budget. We find that introducing interventions earlier generally yields more robust models with no increase in overrefusal rates, with the clearest benefits appearing after downstream, benign finetuning. We also see clear benefits in the steerability of models towards safer generations. Finally, we observe that earlier interventions reshape internal representations: linear probes more cleanly separate safe vs harmful examples. Overall, these results argue for incorporating safety signals early in pretraining, producing models that are more robust to downstream finetuning and jailbreaking, and more reliable under both standard and safety-aware inference procedures.
@@ 1 Introduction
As such, recent work has explored removing all harmful content from pretraining data via filtering approaches (O’Brien et al., 2025). Other work, however, has demonstrated that this could lead to a lack of understanding of harmful content, making models more prone to comply with harmful requests (Li et al., 2025). One recent work of interest (Maini et al., 2025) has explored various synthetic pretraining interventions to (i) maintain information and content from harmful pretraining examples, while (ii) removing explicitly harmful content or providing sufficient context to understand why particular pretraining data may be harmful. Safety pretraining interventions show promise for safer language model behavior, in terms of refusal to comply with harmful queries, adversarial jailbreaks, and even after benign finetuning.
Despite this progress, a central question remains underexplored: when during pretraining should safety interventions be introduced? Curriculum learning (Bengio et al., 2009; Graves et al., 2017) suggests that the order of examples (in terms of topic or difficulty) can qualitatively change downstream performance. However, such topics have been demonstrated primarily in non-language domains, with some work studying the impact of data mixture components at a more coarse-grained level (Xie et al., 2023; Jiang et al., 2024; Ye et al., 2025) or other work introducing benefits of different stages during pretraining (Blakeney et al., 2024). Intuitively, safety interventions introduced only at the end may be too superficial and be easily washed away by downstream finetuning applications, while signals introduced too early or too strongly may degrade standard capabilities and helpfulness or may over-regularize the model.
In this paper, we empirically study the impacts of introducing safety pretraining data at different times during pretraining. We adopt the various pretraining interventions from the work of Maini et al. (2025), such as synthetic rephrasing and contextualization, metadata annotation, and refusal training, and introduce them at different stages (i.e., after 0%, 20%, or 60% of the total amount of 600B tokens) of pretraining. All of our experiments revolve around the experimental setting of pretraining a 1.7B parameter language model with a SmolLM2 architecture (Allal et al., 2025) on a pretraining dataset containing FineWeb-Edu (Penedo et al., 2024), math, code, and other synthetic data (Allal et al., 2025). To evaluate the impact of timing of these interventions, we analyze the safety of the resulting models along multiple axes (i.e., base models, instruction-tuned models, and models after finetuning) and analyze how well the models separate harmful and safe content in their representation space.
@@
We generally find that introducing such interventions earlier leads to greater robustness of base models (i.e., before any instruction tuning) and to greater robustness of instruction-tuned models after a round of benign finetuning. Particularly, we see larger impacts when using a modified inference-time algorithm for safer generations (i.e., a safety-focused beam search). This is complemented by our representation analysis, where we explore the impact of the timing of such pretraining data on the notion of safe versus unsafe data in representation space. In those results, we also find that introducing such topics earlier leads to a greater distinction between harmful and safe content in the representation space, translating to a linear probe yielding higher performance at separating these classes in the last layer representation space of the trained language models. Finally, we find that introducing safety interventions early (i.e., at 20% or 60%) during pretraining leads to significantly more robustness under a threat model of an adversarial jailbreak learned via GCG (Zou et al., 2023), when compared to no safety interventions or safety interventions introduced at the beginning of training (i.e., 0%). Overall, our work suggests the incorporation of safety pretraining interventions earlier during training, yielding improvements in robustness to jailbreaks and to harmful user queries after benign finetuning.
@@ 2 Related Work
Most language model pretraining employs simple heuristic filters that remove a small amount of harmful content, such as from particular URLs or explicit profanity (Raffel et al., 2020; Xue, 2020). One early work studied the role of human preferences in language model pretraining (Korbak et al., 2023), finding that it yields lower toxicity scores relative to finetuned models (on the 100M parameter scale). Other works have taken a more explicit data curation approach. More recently, works have studied heuristic filtering based on educational content (Gunasekar et al., 2023; Penedo et al., 2024), which also correlates with (lower levels of) toxicity. Yet, Vidgen et al. (2024) note that harmful examples remain even in these educational datasets. As such, some work explores the impacts of filtering unsafe content from the pretraining corpus, training only on safe examples (Li et al., 2025; O’Brien et al., 2025).
Most recently and relatedly, the work of Maini et al. (2025) studies the role of synthetic data generation to improve levels of safety in pretraining examples, while still retaining as much information as possible, finding large improvements in robustness after a round of benign finetuning. We expand upon this work to study the question of when such interventions should be deployed during pretraining.
@@
Curriculum learning studies how the order and composition of training data shape generalization. Early work argues for presenting examples in a progression (e.g., from simpler to harder), and for adapting the sampling policy as training proceeds (Bengio et al., 2009; Graves et al., 2017). Such approaches have led to success in various settings, such as question-answering (Sachan and Xing, 2016) or reinforcement learning (Zhang et al., 2020). Other work has explored when and what types of certain curriculum learning strategies work, demonstrated on small-scale image classification tasks (Wu et al., 2021). Progress in the language model pretraining realm has been relatively static, likely due to the difficulty in measuring the hardness of content (Abdin et al., 2024), and perhaps even similarity between content (Sam et al., 2025a), as well as the computational demands of pretraining.
@@
In large-scale language pretraining, most practical curricula have focused on adjusting mixture weights over data sources rather than token-level difficulty, largely due to computational complexity. For example, optimizing source mixtures during pretraining can yield gains in downstream performance and stability (Xie et al., 2023), and a line of work explores adaptive sampling or stage-specific mixtures that emphasize more useful distributions as training evolves (Jiang et al., 2024; Ye et al., 2025). Complementary results in a more coarse-grained setting suggest that pretraining in distinct stages can change what models retain and how they transfer (Blakeney et al., 2024). Our study is aligned with this perspective but targets a safety-specific question: instead of asking which domains to upweight, we vary when safety-augmented data enters pretraining. This isolates timing effects and lets us measure whether safety signals introduced earlier versus later have a greater impact.
@@ 3.2 Safety Curricula
A central question of this work is how the timing of safety interventions during pretraining shapes downstream robustness. Prior work on safety pretraining (Maini et al., 2025) has introduced interventions from the very start of training—effectively modifying the full corpus before pretraining begins. Other work has demonstrated that orderings of certain concepts can improve the speed of learning (Chen et al., 2023). Similarly, in the human learning process, we often first learn about general knowledge, such as language and arithmetic, before addressing potentially harmful or controversial content, such as wars and tragedies. At the other extreme, filtering approaches (O’Brien et al., 2025) that permanently and completely remove all harmful content during pretraining potentially lose too many capabilities from the underlying model. We adopt a similar approach here, where we study a curriculum of first training on safe-only data before introducing safety interventions after a certain percentage of tokens of the full 600B token pretraining budget.
To study the role of the timing of introducing safety interventions, we study 4 main settings. The first setting (0%) applies interventions from the very beginning, which directly mirrors the setup in prior Safety Pretraining work (Maini et al., 2025). This captures no pretraining phase on safe-only data. The second setting (20%) delays interventions until after the model has already seen a substantial portion of “safe” general web-scale data, allowing it to acquire broad linguistic and factual knowledge before being guided by safety signals. The third setting (60%) pushes this even further, introducing interventions only in the later stages of training, when much of the model’s base distribution of language has already been established. Finally, the last setting (100%) corresponds to training completely on safe-only data, which is effectively equivalent to filtering unsafe content for the entirety of pretraining. This has been explored in prior work (Maini et al., 2025), as well as in more harm-specific settings such as biological risks (O’Brien et al., 2025).
To also vary the types of interventions employed, we explore the two primary settings from Safety Pretraining (Maini et al., 2025): (i) we first study using all of the interventions in conjunction, including the metadata tagging and inference-time modifications, (ii) using only rephrasing and refusal training without any metadata tagging. The second set of experiments here considers a slightly less aggressive modification to pretraining, only intervening in terms of data curation.
@@ 4.1 Safety Evaluations
4.2 Helpfulness Evaluations
Whenever introducing safety interventions (that lead to potentially increased refusal rates), it is crucial to also assess any impacts on the helpfulness of the resulting models. We evaluate the performance of our models on complying with benign requests from the Alpaca dataset (Taori et al., 2023), using GPT-4o-mini as a judge to evaluate if responses are (1) compliant and helpful, (2) compliant and unhelpful, (3) partial refusal, or (4) full refusal. Desirable responses are categories (1) and (2), representing model compliance with varying levels of helpfulness. Categories (3) and (4) denote overrefusal in this setting.
When evaluating models that use the modified SafeBeam inference algorithm, we also add in a category (5) shortcircuit, which stops generation when the harmful metadata token is produced. This category reflects a separate and practical setting where we can post-process model responses that contain the harmful token at production time, replacing them with a default refusal message taken from a set of known safe completions.
@@ 5.1 Safety Results
Introducing interventions earlier in pretraining consistently improves the safety of base models (Figure 1). We find that introducing safety pretraining interventions during the middle of training (60% or 20%) leads to the safest base models in terms of unsafe completion rate with standard top-k sampling. Not introducing any safety pretraining interventions and training on safe-only data (100%) leads to the most unsafe performance. Our results also suggest that an earlier inclusion of metadata (at 0%) leads to measurable gains when using SafeBeam inference.
@@
A similar pattern holds with the models that have undergone instruction tuning (Figure 3). Models trained with interventions during the middle of pretraining exhibit stronger safety performance than those with no interventions. Again, SafeBeam benefits the most when metadata is incorporated from the beginning (i.e., 0%).
@@
Taken together, our evaluations suggest two main takeaways. First, when using modified SafeBeam inference, introducing metadata-annotated unsafe data earliest in pretraining leads to the best steerability towards safer completions. Secondly, with top-k inference, introducing safety pretraining interventions during the middle of training (either 20% or 60%) leads to the most robust performance. This suggests that allowing the model to first absorb general-purpose safe web-scale data before applying any safety interventions may produce more durable alignment when subsequent benign finetuning is performed.
@@ 5.2 Overrefusal Results
In assessing the impact of the pretraining interventions on the helpfulness of the models, we find that when using standard top-k sampling (Figure 4), there is a slight improvement when incorporating interventions earlier during pretraining (e.g., at 0% or 20%). This is likely due to models being able to learn more efficiently from high-quality synthetic data (Maini et al., 2024). Overall, this suggests that safety pretraining interventions observe no (or only minor) tradeoffs in overrefusal.
When using the SafeBeam algorithm (Figure 5), we find that incorporating pretraining interventions earlier leads to slightly higher overrefusal rates; however, within the compliant responses, helpfulness increases. Therefore, while overall refusals do increase, the quality of compliant responses improves, suggesting improved performance on general instruction-following tasks. A likely hypothesis for this behavior is that during SafeBeam, the most potentially harmful beams are discarded, and the remaining beams are selected to generate model responses. When pretraining interventions are included earlier during training, the model likely better learns the correspondence between the harmful metadata tag and unsafe content, as well as safer generations through the safe synthetic pretraining data. This leads to the remaining beams in SafeBeam being of higher quality and, ultimately, leading to more helpful responses.
@@ 5.4 Representation Results
6 Conclusion
Our work studies the role of different safety curricula, or how the timing of safety pretraining interventions (e.g., when and how safety-augmented data enters pretraining) impacts the robustness of our models. Our experiments show that incorporating safety pretraining interventions indeed help, and the clearest result is that there is much improved robustness after benign finetuning when pretraining interventions are introduced earlier (e.g., at 0% or 20% of the pretraining tokens). This also manifests into impacts on the model’s underlying representation geometry; incorporating interventions and metadata earlier in pretraining leads to greater separation of safe vs unsafe content. Taken together, these findings highlight how the timing of safety interventions is an important choice in shaping a model’s generations and internal representations, offering guidance for building more trustworthy and socially responsible language models.

### v2

@@ Abstract
Prior work has shown that safety interventions applied during pretraining, such as removing and rephrasing harmful content, can substantially improve the robustness of the resulting models. In this paper, we study the fundamental question that prior work has overlooked: “When during pretraining should safety interventions be introduced?” We keep the underlying data sources and pretraining interventions fixed, varying the intervention start time (after 0%, 20%, or 60% of pretraining tokens). We find that the optimal start time is not one-size-fits-all: with standard top-k decoding, introducing interventions after a short initial phase of safe-only pretraining (20%-60%) often yields the strongest robustness, with the clearest benefits emerging after downstream, benign finetuning. In contrast, for safety-aware inference, interventions starting from the beginning improve steerability towards safer generations. Finally, we observe that earlier interventions reshape internal representations: linear probes more cleanly separate safe vs harmful examples. Our results are the first to establish intervention timing as a key curriculum design choice for safety.
@@ 1 Introduction
These vulnerabilities have motivated a shift toward integrating safety directly into pretraining. Recent work has explored removing all harmful content from pretraining data via filtering approaches (O’Brien et al., 2025). Other work, however, has demonstrated that this could lead to a lack of understanding of harmful content, making models more prone to comply with harmful requests (Li et al., 2025). More promisingly, Maini et al. (2025) explored synthetic pretraining interventions that maintain information from harmful examples while removing explicitly harmful content or providing sufficient context about why such content is harmful. These safety pretraining interventions show promise for improving refusal behavior, robustness to adversarial jailbreaks, and resilience to benign finetuning.
Despite this progress, all prior work on safety pretraining applies interventions uniformly from the start of training, implicitly assuming that when safety signals are introduced does not matter. We challenge this assumption. Drawing on insights from curriculum learning (Bengio et al., 2009; Graves et al., 2017), which has shown that the ordering of training examples can qualitatively change downstream performance, we ask: when during pretraining should safety interventions be introduced? This question has natural intuitions on both sides: interventions introduced too late may be superficial and easily removed by downstream finetuning, while interventions introduced too early may degrade capabilities or over-regularize the model before it has learned basic linguistic structure. To our knowledge, we are the first to systematically study this question.
In this paper, we empirically study the impacts of introducing safety interventions at different times during pretraining. We adopt the various pretraining interventions from the work of Maini et al. (2025), such as synthetic rephrasing and contextualization, metadata annotation, and refusal training, and introduce them at different stages (i.e., after 0%, 20%, or 60% of the total amount of 600B tokens) of pretraining. This leads to various amounts of safety interventions token observed during pretraining. Our experiments are run in the setting of pretraining a 1.7B parameter language model with the SmolLM2 architecture (Allal et al., 2025), on a pretraining mixture containing FineWeb-Edu (Penedo et al., 2024), math, code, and other synthetic datasets (Allal et al., 2025). To evaluate how intervention timing affects safety, we analyze the resulting models at multiple stages (base models, instruction-tuned models, and models after benign finetuning). We also study two common deployment choices: (a) standard top-k decoding, and (b) safety-aware inference enabled by metadata. These settings prioritize different properties (robustness vs steerability), so the best intervention timing differs.
@@
We find that the most effective start time depends on the evaluation setting and inference procedure. Under standard top-k decoding, introducing interventions after some initial safe-only pretraining (after 20%–60%) often yields the safest base and instruction-tuned models, and provides the strongest robustness after benign finetuning. In contrast, when using a safety-aware inference-time algorithm (SafeBeam), starting interventions from the beginning yields the most steerability toward safer generations. This is complemented by our representation analysis: introducing interventions earlier makes safe vs harmful examples more linearly separable in the final-layer representation space, especially when metadata is included. Finally, we find that intervening after an initial safe-only phase (20% or 60%) improves robustness to adversarial jailbreaks learned via GCG (Zou et al., 2023), when compared to performing fully safe-only pretraining (100%) and safety interventions applied from the start (0%). Overall, our results highlight curriculum timing as a key design choice: early starts improve steerability under safety-aware inference, while delayed starts can improve robustness under standard decoding and post-benign finetuning evaluation.
@@ 2 Related Work
Most language model pretraining employs simple heuristic filters that remove a small amount of harmful content, such as from particular URLs or explicit profanity (Raffel et al., 2020; Xue, 2020). One early work studied the role of human preferences in language model pretraining (Korbak et al., 2023), finding that it yields lower toxicity scores relative to finetuned models (on the 100M parameter scale). Other works have taken a more explicit data curation approach. More recently, works have studied heuristic filtering based on educational content (Gunasekar et al., 2023; Penedo et al., 2024), which also correlates with (lower levels of) toxicity. Yet, Vidgen et al. (2024) note that harmful examples remain even in these educational datasets.
As such, some work explores the impacts of filtering unsafe content from the pretraining corpus, training only on safe examples (Li et al., 2025; O’Brien et al., 2025). Most recently and relatedly, the work of Maini et al. (2025) studies the role of synthetic data generation to improve levels of safety in pretraining examples, while still retaining as much information as possible, finding large improvements in robustness after a round of benign finetuning. We expand upon this work to study the question of when such interventions should be deployed during pretraining.
@@
Curriculum learning studies how the order and composition of training data shape generalization. Early work argues for presenting examples in a progression (e.g., from simpler to harder), and for adapting the sampling policy as training proceeds (Bengio et al., 2009; Graves et al., 2017). Such approaches have led to success in various settings, such as question-answering (Sachan and Xing, 2016) or reinforcement learning (Zhang et al., 2020). Other work has explored when and what types of certain curriculum learning strategies work, demonstrated on small-scale image classification tasks (Wu et al., 2021). Progress in the language model pretraining realm has been relatively static, likely due to the difficulty in measuring the hardness (Abdin et al., 2024) and similarity of content (Sam et al., 2025a), as well as the computational demands.
@@
In large-scale language pretraining, most practical curricula have focused on adjusting mixture weights over data sources rather than token-level difficulty, largely due to computational complexity. For example, optimizing source mixtures during pretraining can yield gains in downstream performance and stability (Xie et al., 2023), and a line of work explores adaptive sampling or stage-specific mixtures that emphasize more useful distributions as training evolves (Jiang et al., 2024; Ye et al., 2025).
Complementary results in a more coarse-grained setting suggest that pretraining in distinct stages can change what models retain and how they transfer (Blakeney et al., 2024). Our study is aligned with this perspective but targets a safety-specific question: instead of asking which domains to upweight, we vary when safety-augmented data enters pretraining. This allows us to measure whether safety signals introduced earlier versus later have a greater impact.
@@ 3.1 Safety Pretraining Interventions
Table 1: Capability evaluation results across intervention start timings. We report accuracy on standard benchmarks: ARC-Challenge, CommonsenseQA (CS-QA), GSM8K, OpenBookQA, PIQA, TriviaQA, and Winogrande. We observe no obvious degradations in capabilities as we include safety pretraining interventions earlier (exposing models to more safety pretraining data); in fact, earlier pretraining interventions seem to lead to higher levels of capabilities.
Intervention Avg. ARC-C CS-QA GSM8K OpenBookQA PIQA TriviaQA Wino.
0% 37.2% 42.9% 19.7% 7.4% 42.6% 76.6% 12.7% 58.8%
20% 36.7% 43.5% 22.9% 6.1% 39.8% 76.0% 10.0% 58.7%
60% 35.6% 41.8% 21.1% 5.5% 41.6% 75.1% 5.8% 58.0%
100% 36.1% 42.9% 21.6% 4.9% 42.8% 75.1% 6.0% 59.4%
@@ 3.2 Safety Curricula
A central question of this work is how the timing of safety interventions during pretraining shapes downstream robustness. Prior work on safety pretraining (Maini et al., 2025) has introduced interventions from the very start of training—effectively modifying the full corpus before pretraining begins. Other work has demonstrated that orderings of certain concepts can improve the speed of learning (Chen et al., 2023). Similarly, in the human learning process, we often first learn about general knowledge, such as language and arithmetic, before addressing potentially harmful or controversial content, such as wars and tragedies. At the other extreme, filtering approaches (O’Brien et al., 2025) that permanently and completely remove all harmful content during pretraining potentially lose too many capabilities from the underlying model. We adopt a similar approach here, where we study a curriculum of first pretraining for a phase on safe-only data, before introducing safety interventions after a certain percentage of tokens of the full 600B token pretraining budget.
To study the role of the timing of introducing safety interventions, we study 4 main settings. The first setting (0%) applies interventions from the very beginning, which directly mirrors the setup in prior Safety Pretraining work (Maini et al., 2025). This captures no pretraining phase on safe-only data. The second setting (20%) delays interventions until after the model has already seen a substantial portion of “safe” general web-scale data, allowing it to acquire broad linguistic and factual knowledge before any safety interventions. The third setting (60%) introduces interventions only in the later stages of training, when much of the model’s base distribution of language has already been established. Finally, the last setting (100%) corresponds to fully training on safe-only data, which is equivalent to prior work that filters out all unsafe content (O’Brien et al., 2025).
To also vary the types of interventions employed, we explore the two primary settings from Safety Pretraining (Maini et al., 2025): (i) we first study using all of the interventions in conjunction, including the metadata tagging and inference-time modifications, (ii) using only rephrasing and refusal training without any metadata tagging. The second set of experiments here considers a slightly less aggressive modification to pretraining, only in terms of data curation.
@@ 4.1 Safety Evaluations
4.2 Capability and Helpfulness Evaluations
When introducing safety interventions that may increase refusal rates, it is crucial to assess any impact on model capabilities and helpfulness. For capabilities, we evaluate on a suite of standard benchmarks spanning general knowledge (ARC-Challenge, TriviaQA), commonsense reasoning (CommonsenseQA, OpenBookQA, PIQA, Winogrande), and mathematical reasoning (GSM8K). Additional experiment details are deferred to Appendix B.
To measure helpfulness, we evaluate model compliance on benign requests from the Alpaca dataset (Taori et al., 2023). We use GPT-4o-mini as a judge to classify each response into one of four categories: (1) compliant and helpful, (2) compliant but unhelpful, (3) partial refusal, or (4) full refusal. Categories (1) and (2) represent desirable compliance behavior, while categories (3) and (4) indicate overrefusal. For models using the SafeBeam algorithm, we additionally introduce a fifth category: (5) shortcircuit, which occurs when generation terminates upon producing the harmful metadata token. In practice, such responses can be post-processed at serving time by substituting a default refusal message from a set of known safe completions.
@@ 5.1 Safety Results
Timing matters even before instruction tuning (Figure 1). Under standard top-k sampling, introducing interventions in the middle of pretraining (20% or 60%) yields the safest base models, while training on safe-only data (100%) is the least safe. This again complements the findings from prior work (Li et al., 2025; Maini et al., 2025). In contrast, with SafeBeam, incorporating interventions and metadata from the start (0%) provides measurable gains.
@@
A similar pattern holds with the models that have undergone instruction tuning (Figure 3). Models trained with interventions during the middle of pretraining exhibit stronger safety performance under top-k inference than those with no interventions. Again, SafeBeam benefits the most when metadata is incorporated from the beginning (i.e., 0%).
@@
Taken together, our evaluations suggest two main takeaways. First, when using modified SafeBeam inference, introducing metadata-annotated unsafe data earliest in pretraining leads to the best steerability towards safer completions. Secondly, with top-k inference, introducing safety pretraining interventions during the middle of training (after 20% or 60%) leads to the most robust performance. This suggests that allowing the model to first absorb general-purpose safe data before applying any safety interventions may produce alignment that is more robust to subsequent benign finetuning.
5.2 Capability Results
We evaluate whether safety pretraining interventions degrade general model capabilities across various standard benchmarks (Table 1). Overall, we find that safety interventions introduce minimal capability trade-offs and, in some cases, may even improve performance. We attribute this to the model’s efficient learning from high-quality synthetic data (Maini et al., 2024). Taken together, these results indicate that safety pretraining interventions do not meaningfully degrade model capabilities; indeed, earlier interventions appear mildly beneficial on average.
@@ 5.3 Overrefusal Results
We assess the impact of pretraining interventions on model helpfulness using benign requests from the Alpaca benchmark (Taori et al., 2023). Under standard top-$k$ sampling (Figure 4), incorporating interventions earlier during pretraining (0% or 20%) yields a slight improvement in compliance rates. These results suggest that safety pretraining interventions introduce minimal trade-offs in overrefusal behavior under standard inference.
When using the SafeBeam algorithm (Figure 9), we observe a more nuanced pattern: earlier interventions produce modestly higher overrefusal rates, yet the quality of compliant responses improves. The figure reveals that earlier pretraining interventions (0% and 20%) increase the fraction of compliant-and-helpful responses among non-refused outputs, even as overall refusal rates rise slightly. We hypothesize that this reflects improved calibration: early interventions strengthen the correspondence between the harmful metadata tag and genuinely unsafe content, enabling SafeBeam to filter more precisely. The surviving beams thus tend to produce higher-quality completions.
@@ 5.5 Representation Results
6 Discussion
Our work studies the role of different safety curricula, or how the timing of safety pretraining interventions (e.g., when and how safety-augmented data enters pretraining) impacts the robustness of our models. Our experiments show that incorporating safety pretraining interventions indeed help, but the best timing depends on the setting: delayed interventions (e.g., 20%–60%) can yield stronger robustness under standard decoding and against adversarial jailbreaks, while early interventions (0%) improve steerability under safety-aware inference (SafeBeam) and yield clearer separation of safe vs unsafe content in representation space. Taken together, these findings highlight how the timing of safety interventions is an important choice in shaping a model’s behavior, offering guidance for building more trustworthy and responsible language models.
Why does timing matter?
While our experiments establish that intervention timing significantly affects outcomes, understanding the underlying mechanisms remains an open question. We offer several hypotheses for future investigation. For standard top-k decoding, delayed interventions (20%-60%) may succeed because the model benefits from first acquiring general linguistic competence and world knowledge before safety signals are introduced. Early safety interventions might compete with or interfere with learning basic representations, potentially causing the model to develop shallow safety heuristics rather than robust understanding. By contrast, when the model has already formed stable representations of language and concepts, safety interventions may integrate more deeply into these existing structures, making them harder to remove through subsequent finetuning.
For SafeBeam inference, early interventions may help because the metadata-based steering mechanism relies on strong, well-calibrated associations between the harmful content tag and genuinely unsafe material. Introducing these associations from the start of training allows them to be reinforced throughout the entire pretraining process, resulting in more reliable probability estimates that SafeBeam can exploit. This interpretation is supported by our representation analysis (Figure 6), which shows that early interventions with metadata yield the cleanest separation between safe and unsafe content in representation space.
A limitation of our study is that the introduction of safety pretraining interventions at different times during pretraining means that certain models have seen more synthetic contextualized data or refusal data. However, we note our main effects are not monotonic in the number of intervention tokens (i.e., “more intervention tokens” do not consistently yield safer models), suggesting that when safety-augmented data is introduced matters beyond total quantity.